# Decentralized Stochastic Subgradient Methods for Nonsmooth Nonconvex Optimization

### v1

@@ 4 Application: Decentralized Stochastic Subgradient Methods
    For each $i\in[d]$, there exist functions $\{f_{i,j}:j\in[N_{i}]\}$ such that for any $i\in[d]$, it holds that
    $f_{i}({\bm{x}})=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}f_{i,j}({\bm{x}}).$
    Moreover, for each $i\in[d]$ and $j\in[N_{i}]$, $f_{i,j}$ is definable and admits $\mathcal{D}_{f_{i,j}}$ as its definable conservative field.
@@
It is worth mentioning that Assumption 4.1 is mild and reasonable in practical scenarios. As demonstrated in Section 2, the class of definable functions is wide enough to enclose a wide range of the loss functions of common nonsmooth neural networks. Furthermore, [8] demonstrates that for the loss function of any neural network constructed by definable blocks, the results yielded by Automatic Differentiation (AD) algorithms are contained within a definable conservative field for the definable loss function. Moreover, for the objective functions $\{f_{i}\}$, we can choose their definable conservative fields as
$\mathcal{D}_{f_{i}}:=\sum_{j=1}^{N_{i}}\mathcal{D}_{f_{i,j}}({\bm{x}}).$ (4.1)
In addition, recall that $f({\bm{x}})=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}})$ as defined in (2.1), we can choose its definable conservative field $\mathcal{D}_{f}$ as
@@
Then by Proposition 2.5, we can conclude that the set $\{{f({\bm{x}})}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$ is a finite subset of $\mathbb{R}$.
@@ 4.1 Decentralized stochastic subgradient descent
In this subsection, we establish the global convergence for decentralized stochastic subgradient descent (DSGD) by showing that it fits into our proposed framework (DSM). The detailed algorithm of DSGD is presented in Algorithm 1. Intuitively, for each $i\in[d]$, the consensus term $\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}$ enforces ${\bm{x}}_{i,k}$ towards the mean over all the agents (i.e., $\frac{1}{d}\sum_{i\in[d]}{\bm{x}}_{i,k}$). Moreover, as ${\bm{d}}_{i,k}$ refers to the stochastic subgradient of $f_{i}$ at ${\bm{x}}_{i,k}$, the term $-\eta_{k}{\bm{d}}_{i,k}$ can be viewed as a single SGD step to minimize $f_{i}$ in the $i$-th agent.
@@
5:         Sample the index $j\in[N_{i}]$ uniformly at random;
6:         Compute ${\bm{d}}_{i,k}\in\mathcal{D}_{f_{i,j}}({\bm{x}}_{i,k})$;
@@
8:         Communicate and update the local variables ${\bm{x}}_{i,k}=\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}{\bm{d}}_{i,k}$;
@@
    ${\bm{W}}\in\mathbb{R}^{d\times d}$ is a mixing matrix.
@@
    There exists a constant $M_{x}>0$ such that $\sup_{k\geq 0}\left\|{\bm{X}}_{k}\right\|\leq M_{x}$.
  3. 3.
@@
    $\lim_{k\to+\infty}\eta_{k}\log(k)=0,\quad\sum_{i=0}^{+\infty}\eta_{k}=+\infty.$ (4.3)
Assumption 4.3(1) is identical to Assumption 3.1(1), while Assumption 4.3(2) assumes the uniform boundedness of the sequence $\{{\bm{X}}_{k}\}=\{[\bm{x}_{1,k},...,\bm{x}_{d,k}]\}$. Moreover, Assumption 4.3(3) illustrates that the step-sizes diminish in the rate of $o(1/\log(k))$, hence enabling great flexibility in choosing the step-sizes $\{\eta_{k}\}$ in practical scenarios. In addition, as demonstrated in [2, Proposition 1.3], Assumption 4.3(3) implies Assumption 3.1(3) almost surely. In the following theorem, we establish the global convergence of Algorithm 1 by showing that Algorithm 1 fits into our proposed framework with $\mathcal{H}=\mathcal{D}_{f}$, $\Psi=f$ and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$.
@@
Suppose Assumption 4.1 and 4.3 hold for Algorithm 1. Then for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-critical point of $f$ and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
@@
${\bm{X}}_{k+1}={\bm{X}}_{k}{\bm{W}}-\eta_{k}{\bm{D}}_{k},\qquad\text{ where }\quad{\bm{D}}_{k}=[{\bm{d}}_{1,k},\ldots,{\bm{d}}_{d,k}].$ (4.4)
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. Then we set ${\bm{Z}}_{k}={\bm{X}}_{k}$, ${\bm{H}}_{k}=\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}={\bm{D}}_{k}-\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$. Then it is easy to verify that the update scheme (4.4) is equivalent to the update scheme (DSM). Now we verify the validity of Assumption 3.1 and Assumption 3.2. Assumption 3.1(1)-(2) directly follows from Assumption 4.3(1)-(2) and local boundedness of $\mathcal{D}_{f_{i}}$ in (4.1). Moreover, Assumption 3.1(3) follows from Assumption 4.3(3) and [2, Proposition 1.3]. In addition, from the definition of $\mathcal{D}_{f_{i}}$, we can conclude that
${\bm{H}}_{k}{\bm{e}}\in\frac{1}{d}\sum_{i=1}^{d}\mathcal{D}_{f_{i}}({\bm{x}}_{i,k})\subseteq\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{D}_{f_{i}}({\bm{x}}_{i,k})\right).$ (4.5)
Therefore, by choosing $\mathcal{H}_{i}=\mathcal{D}_{f_{i}}$, we verify the validity of Assumption 3.1(4). Furthermore, according to [8, Section 6] and the path-differentiability of $f$, it follows that $f$ is the Lyapunov function of the differential inclusion
$\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-\mathcal{D}_{f}({\bm{x}}),$ (4.6)
and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. In addition, Remark 4.2 illustrates that $\{f({\bm{x}}):{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we verify the validity of Assumption 3.2 with $\mathcal{H}=\mathcal{D}_{f}$, $\Psi=f$, and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. Finally, from Theorem 3.8, we can conclude that almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-stationary point of $f$ and the sequence of function values $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof. ∎
@@ 4.2 Decentralized generalized SGD with momentum
Momentum is a technique to accelerate the subgradient descent method in the relevant direction and dampen oscillations. In the single-agent setting, [46] investigates the global convergence properties of generalized SGD with momentum, where an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is introduced to normalize the update directions of the primal variables. However, to the best of our knowledge, there is no existing work on developing its decentralized version. In this subsection, motivated by the generalized SGD with momentum analyzed in [46], we first introduce the decentralized generalized SGD with momentum (DGSGDm) and present the details in Algorithm 2. Different from DSGD in Algorithm 1, Algorithm 2 introduces the auxiliary variables $\{{\bm{y}}_{i,k}:k\in\mathbb{N}\}$ to track the first-order moment of the stochastic subgradients of $f_{i}$. Therefore, Algorithm 2 requires two rounds of communication to perform local consensus to $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$ and $\{{\bm{y}}_{1,k},\ldots,{\bm{y}}_{d,k}\}$, respectively. Moreover, similar to the generalized SGD analyzed in [46], Algorithm 2 introduces an auxiliary function $\phi$ to normalize the update directions of its primal variables $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$.
@@
The auxiliary function $\phi$ in Algorithm 2 determines how the update directions of the variables $\{{\bm{x}}_{k}\}$ are regularized. Moreover, Algorithm 2 yields different variants of DSGD-type methods with different choices of the auxiliary function $\phi$. For example, Algorithm 2 corresponds to DSGDm when choosing $\phi({\bm{y}})=\frac{1}{2}\left\|{\bm{y}}\right\|^{2}$. Furthermore, when we select $\phi({\bm{y}})=\left\|{\bm{y}}\right\|_{1}$, Algorithm 2 employs the sign map to regularize the update directions of the variables $\{{\bm{x}}_{k}\}$. Under such selection of $\phi$, we can view this algorithm as a decentralized variant of SignSGD analyzed in [3], and denoted it as DSignSGD. Interested readers could refer to [46] for more details on the choices of the auxiliary function $\phi$.
@@
6:         Communicate and update local variable ${\bm{x}}_{i,k+1}\in\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}\partial\phi({\bm{y}}_{i,k})$;
7:         Sample the index $j\in[N_{i}]$ uniformly at random;
8:         Compute ${\bm{d}}_{i,k+1}\in\mathcal{D}_{f_{i,j}}({\bm{x}}_{i,k+1})$;
@@
Suppose Assumption 4.1 and 4.3 hold for Algorithm 2. Moreover, we assume that the auxiliary function $\phi$ is convex and admits a unique minimizer at ${\bm{0}}$. Then for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-stationary point of $f$ and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
@@
${\bm{X}}_{k+1}\in{}$ ${\bm{X}}_{k}{\bm{W}}-\eta_{k}{\bm{T}}_{k},\quad{\bm{T}}_{k}\in[\partial\phi({\bm{y}}_{1,k}),\ldots,\partial\phi({\bm{y}}_{d,k})],$ (4.7)
@@
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. Then by setting ${\bm{Z}}_{k}=\left[\begin{smallmatrix}{\bm{X}}_{k}\\ {\bm{Y}}_{k}\end{smallmatrix}\right]$, ${\bm{H}}_{k}=\mathbb{E}\left[\left[\begin{smallmatrix}{\bm{T}}_{k}\\ \tau{\bm{Y}}_{k}{\bm{W}}-\tau{\bm{D}}_{k+1}\end{smallmatrix}\right]|\mathcal{F}_{k}\right]$, and $\Xi_{k+1}=\left[\begin{smallmatrix}{\bm{T}}_{k}\\ \tau{\bm{Y}}_{k}{\bm{W}}-\tau{\bm{D}}_{k+1}\end{smallmatrix}\right]-{\bm{H}}_{k}$, we can conclude that the update scheme (4.7) can be reshaped as
${\bm{Z}}_{k+1}={\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}).$ (4.8)
@@
Now we verify the validity of Assumption 3.1 and Assumption 3.2. Assumption 3.1(1) is identical to Assumption 4.3(1). Moreover, from the uniform boundedness of $\{{\bm{X}}_{k}\}$ in Assumption 4.3(1) and the local boundedness of $\{\mathcal{D}_{f_{i}}\}$, we can conclude that $\{{\bm{D}}_{k}\}$ is uniformly bounded almost surely. Together with the update scheme of ${\bm{Y}}_{k}$, we can conclude that $\left\|{\bm{Y}}_{k+1}\right\|\leq(1-\tau\eta_{k})\left\|{\bm{Y}}_{k}\right\|+\tau\eta_{k}\left\|{\bm{D}}_{k+1}\right\|$, hence there exists a constant $M_{Y}>0$ such that $\sup_{k\geq 0}\left\|{\bm{Y}}_{k}\right\|\leq\max\{\left\|{\bm{Y}}_{0}\right\|,\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\}<M_{Y}$. Then combining with local boundedness of $\partial\phi$, it can be referred that $\{{\bm{T}}_{k}\}$ is uniformly bounded almost surely. This illustrates that Assumption 3.1(2) holds. Moreover, by the definition of $\{\Xi_{k}\}$, it is straightforward to see that $\{\Xi_{k}\}$ is uniformly bounded almost surely. Then Assumption 4.3(3) and [2, Remark 1.5] illustrates that Assumption 3.1(3) holds almost surely for Algorithm 2. In addition, from the definition of $\mathcal{D}_{f_{i}}$, let $\delta_{k}=\left\|{\bm{X}}_{k+1}-{\bm{X}}_{k}\right\|_{\mathrm{F}}$ and
$\mathcal{H}_{i}({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\mathcal{D}_{f_{i}}({\bm{x}})\end{smallmatrix}\right],$ (4.9)
we can conclude that
${\bm{H}}_{k}{\bm{e}}\in\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}^{\delta_{k}}_{i}({\bm{x}}_{i,k},{\bm{y}}_{i,k}).$ (4.10)
Followed by the same techniques in Lemma 3.4 and Proposition 3.5, we can prove that $\delta_{k}\to 0$. This verifies the validity of Assumption 3.1(4). Furthermore, from the definition of $f$ in Assumption 4.1(1), we attain that
$\mathcal{H}({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\mathcal{D}_{f}({\bm{x}})\end{smallmatrix}\right].$ (4.11)
Then together with [46, Proposition 4.5] and the path-differentiability of $f$, we can conclude that $\Psi({\bm{x}},{\bm{y}})=f({\bm{x}})+\frac{1}{\tau}\phi({\bm{y}})$ is the Lyapunov function of the differential inclusion
$\left(\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t},\frac{\mathrm{d}{\bm{y}}}{\mathrm{d}t}\right)\in-\mathcal{H}({\bm{x}},{\bm{y}}),$ (4.12)
@@
and admits $\{({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ as its stable set. Together with the definability of $f$ and $\mathcal{D}_{f}$, we can conclude that the set $\{\Psi({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ is finite in $\mathbb{R}$. Therefore, Assumption 3.2 holds for Algorithm 2. Finally, from Theorem 3.8, we can conclude that almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-stationary point of $f$ and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof. ∎

### v2

@@ 4 Application: Decentralized Stochastic Subgradient-based Methods
    For each $i\in[d]$ and ${\bm{s}}_{l}\in\mathcal{S}_{i}$, $F(\cdot;{\bm{s}}_{l})$ is definable and admits a definable conservative field $D_{F(\cdot;{\bm{s}}_{l})}$.
@@
It is worth mentioning that Assumption 4.1 is mild and reasonable in practical scenarios. As demonstrated in [29, 28], the class of definable functions is wide enough to enclose a wide range of the loss functions of common nonsmooth neural networks. Moreover, [28] demonstrates that there exists a definable conservative field for any definable loss function constructed by definable blocks, within which the results yielded by Automatic Differentiation (AD) algorithms are contained.
Under Assumption 4.1 (1), we can choose the definable conservative field of $f_{i}$ as
$\mathcal{D}_{f_{i}}({\bm{x}}):=\frac{1}{|\mathcal{S}_{i}|}\sum_{l=1}^{|\mathcal{S}_{i}|}D_{F(\cdot;{\bm{s}}_{l})}({\bm{x}}).$ (4.1)
Recall that $f({\bm{x}})=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}})$ as defined in (1.1), we can choose its definable conservative field $\mathcal{D}_{f}$ as
@@
Then by [28, Theorem 5], we conclude that the set $\{{f({\bm{x}})}:{\bm{0}}\in D_{f}({\bm{x}})\}$ is a finite subset of $\mathbb{R}$.
@@ 4.1 Decentralized stochastic subgradient descent
In this subsection, we establish the global convergence for decentralized stochastic subgradient descent (DSGD) by showing that it fits into our proposed framework (DSM). The detailed algorithm of DSGD is presented in Algorithm 1. Intuitively, for each $i\in[d]$, the consensus term $\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}$ enforces ${\bm{x}}_{i,k}$ towards the mean over all the agents (i.e., $\frac{1}{d}\sum_{i\in[d]}{\bm{x}}_{i,k}$). Moreover, as ${\bm{d}}_{i,k}$ refers to the (mini-batch) stochastic subgradient of $f_{i}$ at ${\bm{x}}_{i,k}$, the term $-\eta_{k}{\bm{d}}_{i,k}$ can be viewed as a single SGD step to minimize $f_{i}$ in the $i$-th agent.
@@
5:         Randomly select a mini-batch $\mathcal{B}_{i,k}\subseteq\mathcal{S}_{i}$;
6:         Compute ${\bm{d}}_{i,k}\in\frac{1}{|\mathcal{B}_{i,k}|}\sum_{l\in\mathcal{B}_{i,k}}D_{F(\cdot;{\bm{s}}_{l})}({\bm{x}})$;
@@
8:         Communicate and update the local variables
${\bm{x}}_{i,k+1}=\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}{\bm{d}}_{i,k};$
@@
    There exists a constant $M_{x}>0$ such that
    $\sup_{k\geq 0}\left\|{\bm{X}}_{k}\right\|\leq M_{x}.$
@@
    $\lim_{k\to+\infty}\eta_{k}\log(k)=0,\text{ and }\sum_{i=0}^{+\infty}\eta_{k}=+\infty.$
Assumption 4.3(1) assumes the uniform boundedness of the sequence $\{{\bm{X}}_{k}\}=\{[\bm{x}_{1,k},...,\bm{x}_{d,k}]\}$. Moreover, Assumption 4.3(2) illustrates that the step-sizes diminish in the rate of $o(1/\log(k))$, hence enabling great flexibility in choosing the step-sizes $\{\eta_{k}\}$ in practical scenarios. In addition, as demonstrated in [30, Proposition 1.3], Assumption 4.3(2) implies Assumption 3.1(2) almost surely.
In the following theorem, we establish the global convergence of Algorithm 1 by showing that Algorithm 1 fits into our proposed framework with $\Phi=\mathcal{D}_{f}$, $\psi=f$ and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$.
@@
Suppose Assumption 4.1 and 4.3 hold for Algorithm 1. Then for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-critical point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
@@
${\bm{X}}_{k+1}={\bm{X}}_{k}{\bm{W}}-\eta_{k}{\bm{D}}_{k},~{}{\bm{D}}_{k}=[{\bm{d}}_{1,k},\ldots,{\bm{d}}_{d,k}].$ (4.3)
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. Then we set ${\bm{Z}}_{k}={\bm{X}}_{k}$, ${\bm{H}}_{k}=\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}={\bm{D}}_{k}-\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$. Then it is easy to verify that the update scheme (4.3) is equivalent to the update scheme (DSM).
Now we verify the validity of Assumption 3.1 and Assumption 3.3. Assumption 3.1(1) directly follows from Assumption 4.3(1) and local boundedness of the conservative field $D_{f_{i}}$ in (4.1). Moreover, Assumption 3.1(2) follows from Assumption 4.3(2) and [30, Proposition 1.3]. In addition, from the fact that $\mathbb{E}[{\bm{d}}_{i,k}|\mathcal{F}_{k}]\in\mathbb{E}_{\mathcal{B}_{i,k}}[D_{\mathcal{B}_{i,k}}({\bm{x}}_{i,k})]=D_{f_{i}}({\bm{x}}_{i,k})$, we can conclude that
$\frac{1}{d}{\bm{H}}_{k}{\bm{1}}_{d}\in\frac{1}{d}\sum_{i=1}^{d}D_{f_{i}}({\bm{x}}_{i,k})\subseteq\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}D_{f_{i}}({\bm{x}}_{i,k})\right).$
Therefore, by choosing $\Phi_{i}=\mathcal{D}_{f_{i}}$, we have verified the validity of Assumption 3.1(3).
Furthermore, according to [28, Section 6] and the path-differentiability of $f$, it follows that $f$ is the Lyapunov function of the differential inclusion $\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-\mathcal{D}_{f}({\bm{x}})$, and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. In addition, Assumption 4.1(1) illustrates that $\{f({\bm{x}}):{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we verify the validity of Assumption 3.3 with $\Phi=\mathcal{D}_{f}$, $\psi=f$, and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$.
Finally, from Theorem 3.9, we can conclude that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-stationary point of (DOP) and the sequence of function values $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof.
@@ 4.2 Decentralized generalized SGD with momentum
Momentum is a technique to accelerate the subgradient descent method in the relevant direction and dampen oscillations. In the single-agent setting, [36] investigates the global convergence properties of generalized SGD with momentum, where an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is introduced to normalize the update directions of the primal variables. However, to the best of our knowledge, there is no existing work on developing its decentralized version.
In this subsection, motivated by the generalized SGD with momentum analyzed in [36], we first introduce the decentralized generalized SGD with momentum (DGSGDm) and present the details in Algorithm 2. Different from DSGD in Algorithm 1, Algorithm 2 introduces the auxiliary variables $\{{\bm{y}}_{i,k}:k\in\mathbb{N}\}$ to track the first-order moment of the stochastic subgradients of $f_{i}$. Therefore, Algorithm 2 requires two rounds of communication to perform local consensus to $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$ and $\{{\bm{y}}_{1,k},\ldots,{\bm{y}}_{d,k}\}$, respectively. Moreover, similar to the generalized SGD analyzed in [36], Algorithm 2 introduces an auxiliary function $\phi$ to normalize the update directions of its primal variables $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$.
@@
The auxiliary function $\phi$ in Algorithm 2 determines how the update directions of the variables $\{{\bm{x}}_{k}\}$ are regularized. Hence, Algorithm 2 yields different variants of DSGD-type methods with different choices of the auxiliary function $\phi$. For example, Algorithm 2 corresponds to DSGD-M when choosing $\phi({\bm{y}})=\frac{1}{2}\left\|{\bm{y}}\right\|^{2}$. Furthermore, when we select $\phi({\bm{y}})=\left\|{\bm{y}}\right\|_{1}$, Algorithm 2 employs the sign map to regularize the update directions of the variables $\{{\bm{x}}_{k}\}$. Under such selection of $\phi$, we can view this algorithm as a decentralized variant of SignSGD analyzed in [43], and denoted it as DSignSGD. Interested readers could refer to [36] for more details on the choices of the auxiliary function $\phi$.
@@
6:         Communicate and update the local variable
${\bm{x}}_{i,k+1}\in\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}\partial\phi({\bm{y}}_{i,k});$
7:         Randomly select a mini-batch $\mathcal{B}_{i,k+1}\subseteq\mathcal{S}_{i}$;
8:         Compute ${\bm{d}}_{i,k+1}\in\frac{1}{|\mathcal{B}_{i,k+1}|}\sum_{l\in\mathcal{B}_{i,k+1}}D_{F(\cdot;{\bm{s}}_{l})}({\bm{x}})$;
@@
Suppose Assumption 4.1 and 4.3 hold for Algorithm 2. Moreover, we assume that the auxiliary function $\phi$ is convex and admits a unique minimizer at ${\bm{0}}$. Then for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-stationary point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
@@
${\bm{X}}_{k+1}\in{}$ ${\bm{X}}_{k}{\bm{W}}-\eta_{k}{\bm{T}}_{k},~{}{\bm{T}}_{k}\in[\partial\phi({\bm{y}}_{1,k}),\ldots,\partial\phi({\bm{y}}_{d,k})],$ (4.4)
@@
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l},~{}{\bm{Y}}_{l}:l\leq k\})$. Then by setting ${\bm{Z}}_{k}=\left[\begin{smallmatrix}{\bm{X}}_{k}\\ {\bm{Y}}_{k}\end{smallmatrix}\right]$, ${\bm{H}}_{k}=\mathbb{E}\left[\left[\begin{smallmatrix}{\bm{T}}_{k}\\ \tau{\bm{Y}}_{k}{\bm{W}}-\tau{\bm{D}}_{k+1}\end{smallmatrix}\right]|\mathcal{F}_{k}\right]$, and $\Xi_{k+1}=\left[\begin{smallmatrix}{\bm{T}}_{k}\\ \tau{\bm{Y}}_{k}{\bm{W}}-\tau{\bm{D}}_{k+1}\end{smallmatrix}\right]-{\bm{H}}_{k}$, we can conclude that the update scheme (4.4) can be reshaped as ${\bm{Z}}_{k+1}={\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})$.
@@
Now we verify the validity of Assumption 3.1 and Assumption 3.3. From the uniform boundedness of $\{{\bm{X}}_{k}\}$ in Assumption 4.3(1) and the local boundedness of $\mathcal{D}_{f_{i}}$, we can conclude that $\{{\bm{D}}_{k}\}$ is uniformly bounded almost surely. Together with the update scheme of ${\bm{Y}}_{k}$, we can conclude that $\left\|{\bm{Y}}_{k+1}\right\|\leq(1-\tau\eta_{k})\left\|{\bm{Y}}_{k}\right\|+\tau\eta_{k}\left\|{\bm{D}}_{k+1}\right\|$, hence there exists a constant $M_{Y}>0$ such that $\sup_{k\geq 0}\left\|{\bm{Y}}_{k}\right\|\leq\max\{\left\|{\bm{Y}}_{0}\right\|,\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\}<M_{Y}$. Then combining with local boundedness of $\partial\phi$, it can be referred that $\{{\bm{T}}_{k}\}$ is uniformly bounded almost surely. This illustrates that Assumption 3.1(1) holds.
Moreover, by the definition of $\{\Xi_{k}\}$, it is straightforward to see that $\{\Xi_{k}\}$ is uniformly bounded almost surely. Then Assumption 4.3(2) and [30, Remark 1.5] illustrates that Assumption 3.1(2) holds almost surely for Algorithm 2. In addition, from the definition of $\mathcal{D}_{f_{i}}$, let $\delta_{k}=\left\|{\bm{X}}_{k+1}-{\bm{X}}_{k}\right\|_{\mathrm{F}}$ and $\Phi_{i}({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\mathcal{D}_{f_{i}}({\bm{x}})\end{smallmatrix}\right]$, we can conclude that $\frac{1}{d}{\bm{H}}_{k}{\bm{1}}_{d}\in\frac{1}{d}\sum_{i=1}^{d}\Phi^{\delta_{k}}_{i}({\bm{x}}_{i,k},{\bm{y}}_{i,k})$. Followed by the same techniques in Lemma 3.5 and Proposition 3.6, we can prove that $\delta_{k}\to 0$. This verifies the validity of Assumption 3.1(3).
Furthermore, from the definition of $f$ in Assumption 4.1(1), we attain that $\Phi({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\mathcal{D}_{f}({\bm{x}})\end{smallmatrix}\right]$. Then together with [36, Proposition 4.5] and the path-differentiability of $f$, we can conclude that $\psi({\bm{x}},{\bm{y}})=f({\bm{x}})+\frac{1}{\tau}\phi({\bm{y}})$ is the Lyapunov function of the differential inclusion $\left(\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t},\frac{\mathrm{d}{\bm{y}}}{\mathrm{d}t}\right)\in-\Phi({\bm{x}},{\bm{y}})$, and admits $\{({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ as its stable set. Together with the definability of $f$ and $\mathcal{D}_{f}$, we can conclude that the set $\{\psi({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ is finite in $\mathbb{R}$. Therefore, Assumption 3.3 holds for Algorithm 2.
@@
Finally, from Theorem 3.9, we can conclude that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-stationary point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof. ∎